5. Snapdragons have incomplete dominance with the RR genotype producing red flowers, the Rr genotype producing pink flowers and the rr genotype producing white flowers. In a population of 10,000 snapdragons, 6500 are red, 2000 are pink, and 1500 are white. What is the frequency of the r allele?

a.

.2

b.

.4

c.

.5

d.

.125

e.

.25
Answer #1

Ans.) Option B : 0.4

Explanation:

According to Hardy-Weinberg principle:

p2 + 2pq + q2 = 1 and p + q = 1

Where,

p2: denotes "frequency of individual with red flower and genotype RR"

2pq : denotes " frequency of individual with pink flower and genotype Rr"

q2: denotes "frequency of individual with white flower and genotype rr"

p : denotes "frequency of dominant R gene in the population"

q : denotes "frequency of recessive gene in the popuation"

Given that:

Total population size = 10,000

No. of individual with red phenotype (RR) = 6500

So, p2 (frequency of individual with red flower and genotype RR) = 6500/ 10,000 = 0.65

No. of individual with pink phenotype (Rr) = 2000

So, 2pq (frequency of individual with pink flower and genotype Rr) = 2000/ 10,000 = 0.20

No. of individual with white flower phenotype (rr) = 1500

q2 (frequency of individual with white flower and genotype rr) = 1500/ 10,000 = 0.15

So, frequency of r allele = p = = = 0.3872 (approx . 0.4)